# Performance of Conformal Prediction in Capturing Aleatoric Uncertainty

### v1

@@ 6 Discussion
We evaluate the common assumption that conformal predictors capture aleatoric uncertainty. We perform this by measuring their correlation and similarity with the ambiguity present in human annotations across four datasets. Our experiments use a diverse set of datasets with varying degrees of class overlap from human annotators. Overall, we found that conformal predictors find it challenging to capture the class overlap recorded using multiple human annotators.
In agreement with previous works [2], we see that the conformal predictors lead to an increased coverage of the true class for all the datasets and models. This is presented in Tables 2-5.
Even though conformal predictors often show a strong correlation with aleatoric uncertainty quantifiers such as the softmax entropy (See Table 7), they usually fail to capture the inherent class overlap in the employed datasets and show substantial differences from human annotations. For the challenging dataset, MLRSNet, where 97.2% of the images have overlapping classes (or received more than two distinct human labels), the Spearman’s $r_{s}$ between the prediction set size and class overlap was always very weak. This highlights the particular shortcomings of prediction sets when presented with challenging tasks. We observe weak and a few moderate correlations in cases where class overlap is low, as in the ImageNet-Real dataset (there is class overlap in only 15.8% of the total images from the ImageNet-ReaL dataset) or where the number of categories is small as in the FER+ dataset (the FER+ has only seven categories, which is the lowest number of categories relative to the rest of the datasets).
@@ Acknowledgments
Appendix
@@
We assess the similarity between the prediction sets of the conformal predictors and human annotations using the metrics precision, recall, subset-accuracy (S. Acc), and hamming loss. The results on all the datasets are presented in Tables A1, A2, A3 and A4.

### v2

@@ 6 Discussion
We evaluate the common assumption that conformal predictors capture aleatoric uncertainty. We perform this by measuring their correlation and similarity with the ambiguity present in human annotations across four datasets. Our experiments use a diverse set of datasets with varying degrees of class overlap from human annotators. Overall, we found that conformal predictors find it challenging to capture the class overlap recorded using multiple human annotators. We also observe similar results when considering entropy in human annotations.
In agreement with previous works [2], we see that the conformal predictors lead to an increased coverage of the true class for all the datasets and models. This is presented in Table 4 and in Tables A1-A4 in the Supplementary material.
Even though conformal predictors often show a strong correlation with aleatoric uncertainty quantifiers such as the softmax entropy (See Table 2), they usually fail to capture the inherent class overlap and entropy in human annotations in the employed datasets, and show substantial differences from human annotations. For the challenging dataset, MLRSNet, where 97.2% of the images have overlapping classes (or received more than two distinct human labels), the Spearman’s $r_{s}$ between the prediction set size and class overlap was always very weak. This highlights the particular shortcomings of prediction sets when presented with challenging tasks. We observe weak and a few moderate correlations in cases where class overlap is low, as in the ImageNet-Real dataset (there is class overlap in only 15.8% of the total images from the ImageNet-ReaL dataset) or where the number of categories is small as in the FER+ dataset (the FER+ has only seven categories, which is the lowest number of categories relative to the rest of the datasets).
In this paper, we focus on aleatoric uncertainty as captured by class overlap. We also quantify aleatoric uncertainty using human annotation entropy whenever such information is available. This form of uncertainty is irreducible because it arises from the nature of the datasets and the label space itself. We use multiple human annotators per instance and define class overlap as the number of distinct labels assigned to that instance; more distinct labels indicate greater intrinsic ambiguity. For example, it is expected that a satellite image of an airport can contain airplanes. So, if we ask several annotators for a single label, we may reasonably get the labels airport and airplane for the same image. Similar patterns occur in subjective tasks such as emotion recognition, where different but defensible choices can yield multiple distinct labels for the same face. By grounding our evaluation in these naturally occurring ambiguities, we can directly assess whether conformal predictors capture the intrinsic uncertainty present in real data.
By contrast, epistemic uncertainty reflects limitations of the model or data and is, in principle, reducible (e.g., labeling error). To minimize any potential leakage from epistemic factors, we (i) use multiple annotators per item with consistent instructions, (ii) assess ECE and make sure that the trained models have low calibration error, and (iii) rely on datasets and training protocols that yield strong performance in prior work and reproduce these results in our experiments.
@@ Acknowledgments
Supplementary material
Paper #1305 : Performance of Conformal Prediction in Capturing Aleatoric Uncertainty
Models’ and conformal predictors’ performance
The models’ accuracy and the conformal predictors’ coverage, SSC, and mean set size ($\bar{w}$) on all four datasets are presented in Table A1 to Table A4.
@@
We assess the similarity between the prediction sets of the conformal predictors and human annotations using the metrics precision, recall, subset-accuracy (S. Acc), and Hamming loss. The results on all the datasets are presented in Tables A5, A6, A7 and A8.
Correlation between conformal prediction sets and human annotation entropy
We present the results of the correlation analysis between prediction sets and human annotation entropy on the CIFAR-10H and FER+ datasets in Table A9.
Expected Calibration Error
Table A10 presents the Expected Calibration Error (ECE) using M=15 bins for all the models.
Table A1: Models’ and conformal predictors’ performance on the CIFAR-10H dataset.
Models Accuracy LAC APS RAPS
Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$
ResNet18 0.930 0.944 0.894 1.040 0.977 0.953 1.357 0.978 0.952 1.361
ResNet34 0.933 0.953 0.833 1.062 0.979 0.953 1.369 0.976 0.950 1.304
ResNet50 0.936 0.953 0.500 1.050 0.974 0.939 1.263 0.975 0.959 1.285
VGG-16 0.941 0.955 0.853 1.049 0.982 0.930 1.328 0.979 0.882 1.263
VGG-19 0.939 0.951 0.000 1.039 0.976 0.907 1.264 0.973 0.895 1.207
Table A2: Models’ and conformal predictors’ performance on the MLRSNet dataset.
Models Accuracy LAC APS RAPS
Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$
ResNet18 0.927 0.951 0.000 1.081 0.984 0.889 1.575 0.985 0.900 1.624
ResNet34 0.923 0.951 0.831 1.090 0.982 0.500 1.500 0.982 0.750 1.526
ResNet50 0.918 0.949 0.780 1.113 0.978 0.667 1.504 0.978 0.667 1.508
VGG-16 0.872 0.950 0.750 1.613 0.955 0.833 1.798 0.963 0.887 2.059
VGG-19 0.844 0.955 0.889 2.106 0.958 0.875 2.352 0.974 0.953 3.386
Table A3: Models’ and conformal predictors’ performance on the FER+ dataset.
Models Accuracy LAC APS RAPS
Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$
ResNet18 0.825 0.941 0.818 1.561 0.961 0.920 2.250 0.960 0.920 2.233
ResNet34 0.839 0.939 0.934 1.484 0.958 0.936 2.085 0.958 0.933 2.138
ResNet50 0.838 0.940 0.857 1.493 0.959 0.800 2.027 0.961 0.857 2.148
VGG-16 0.835 0.946 0.875 1.530 0.970 0.952 2.274 0.970 0.951 2.318
VGG-19 0.823 0.943 0.935 1.561 0.964 0.875 2.222 0.966 0.923 2.316
Table A4: Models’ and conformal predictors’ performance on the ImageNet-ReaL dataset.
Models Accuracy LAC APS RAPS
Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$
ResNet34 0.759 0.899 0.681 2.157 0.940 0.000 8.894 0.940 0.500 8.943
ResNet50 0.785 0.903 0.000 1.760 0.946 0.000 7.768 0.946 0.000 7.892
VGG-16 0.745 0.902 0.000 2.461 0.943 0.000 10.690 0.943 0.000 10.587
VGG-19 0.751 0.903 0.000 2.355 0.946 0.000 10.764 0.946 0.000 10.692
@@
Table A9: Spearman’s rank correlation coefficient, $r_{s}$, $p<.001$, between conformal prediction set sizes and human annotation entropy.
Models CIFAR-10H FER+
LAC APS RAPS LAC APS RAPS
ResNet18 0.140 0.255 0.255 0.404 0.354 0.365
ResNet34 0.142 0.233 0.228 0.413 0.272 0.272
ResNet50 0.160 0.246 0.240 0.413 0.309 0.310
VGG-16 0.153 0.254 0.246 0.457 0.297 0.301
VGG-19 0.162 0.240 0.241 0.429 0.296 0.276
DenseNet121 0.025 0.230 0.228 0.437 0.191 0.191
DenseNet161 0.058 0.228 0.220 0.454 0.277 0.272
MobileNet-v2 0.124 0.269 0.268 0.427 0.337 0.343
Table A10: Expected Calibration Error (ECE) with M = 15 bins.
Models CIFAR-10H MLRSNet FER+ ImageNet-ReaL
ResNet18 0.021 0.013 0.032 0.021
ResNet34 0.026 0.023 0.035 0.027
ResNet50 0.023 0.030 0.026 0.032
VGG-16 0.016 0.100 0.037 0.020
VGG-19 0.022 0.119 0.053 0.020
DenseNet121 0.020 0.018 0.034 0.020
DenseNet161 0.021 0.019 0.038 0.047
MobileNet-v2 0.025 0.012 0.027 0.020